OpenStudy (anonymous):

A disc jockey must allow time for 24 minutes of commercials every hour, along with 4 minutes for news, 3 minutes for weather, and 2 minutes for public-service announcements. If each record lasts and average of 3 minutes, how many records per hour can the DJ play?

OpenStudy (anonymous):

First find the total amount of time the DJ has to play records. To do this, take his initial time (60) and subtract the time needed for commercials, news, weather, and public service announcements.

OpenStudy (anonymous):

these are the choices. 3 seconds per hour 6 seconds per hour 9 seconds per hour and 12 seconds per hour

OpenStudy (anonymous):

Okay, but let's just work on solving the problem. So can you subtract the time needed for commercials, news, weather, and public service announcements from the total time?

OpenStudy (anonymous):

so i take 60, and subtract the commericals, news, weather and public service announcements ?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

would it be 12 seconds per hour then?

OpenStudy (anonymous):

Well what did you get when you did the subtraction?

OpenStudy (anonymous):

well, i got 51? but this doesnt make sense to me :/

OpenStudy (anonymous):

So we have 60. Now we subtract 24. So 60-24=?

OpenStudy (anonymous):

36 !

OpenStudy (anonymous):

Good, so now what is 36-4=?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

And now 32-3=?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

And finally, 29-2=?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

Ok, nice. So now we know that the DJ has 27 minutes not spent playing records. To find out how many records he can play, we'll use the formula \[\large \sf \frac{Time~the~DJ~can~play~records}{Time~per~records}~=~Amount~of~records~played\]

OpenStudy (anonymous):

Whoops, I had a typo. I meant to say SPENT playing record. So imagine the "not" doesn't exist.

OpenStudy (anonymous):

okay, so now how do i do this.

OpenStudy (anonymous):

Well all we have to do is plug in the numbers. How much time does the DJ have to play records? (Hint: We just subtracted to find this out). And how much time does a record take? (Hint: It's in the original question)

OpenStudy (anonymous):

is it 27- 24?

OpenStudy (anonymous):

No, we are dividing.

OpenStudy (anonymous):

Can you answer the questions I asked in my previous post?

OpenStudy (anonymous):

im trying D: okay so it will be 27 divided by 3? im so confused with plugging in these numbers.

OpenStudy (anonymous):

Yes, it would be 27 divided by 3 :).

OpenStudy (anonymous):

thank you so much !!!

OpenStudy (anonymous):

No problem :)